Session: May 1998 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Maria J. Rioja at Sat Jun 6 00:45:05 METDST 1998) Success
No known problems.

Torun (K. Borkowski at Sat Jun 6 14:09:12 METDST 1998) Failure
The recorder is still unusable. The problem has been traced to the headstack movement mechanism.

Onsala (John Conway at Sun Jun 7 17:56:55 METDST 1998) Success
No known problems

Westerbork (Tony Foley at Mon Jun 8 09:56:10 METDST 1998) Success
No known problems

Noto (C.Trigilio at Mon Jun 8 18:04:37 METDST 1998) Minor failures
The operator reports a problem of the time of the FS at 19:33 that caused the lost of scan from 19:34 to 19:40; scan 19:42-19:55 probabily lost.

Jodrell (Alastair Gunn at Tue Jun 9 03:04:35 METDST 1998) Minor failures
Head position did not move between scans 1 to 4, okay from scan 5 (restarted schedule at 13:06 UT). Problem with IFD procedure - may be set to nor,nor instead of nor,alt for some periods prior to fix at about 16:30 UT. Strong RFI at 1673 and 1646 MHz, VC9 very bad.

Cambridge (Alastair Gunn at Tue Jun 9 03:05:19 METDST 1998) Minor failures
chedule started but telescope stowed to fix receiver until on source at 12:07 UT. Low signal as expected in BBCs 7 and 8. Tape lost vacuum at 17:12 and 19:42 but restarted.

Calibration (Jean\-Francois Desmurs at Wed Jun 10 15:58:50 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name em024a.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Fri Jun 12 14:54:23 METDST 1998) Minor failures
could not see 0804 on x-scans, strange ! but 3C196 seen OK.